# MedicoSAM: Towards foundation models for medical image segmentation

### v1

@@ Acknowledgment
The work of Anwai Archit was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - PA 4341/2-1. The work of Luca Freckmann was funded by the DFG under Germany’s Excellence Strategy - EXC 2067/1-390729940. This work was also supported by the Google Research Scholarship “Vision Foundation Models for Bioimage Segmentation”. We gratefully acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Emmy at NHR@Göttingen as part of the NHR infrastructure, under the project nim00007. We would like to thank Sebastian von Haaren for suggestions on improving data visualizations.
[…]
Appendix
We show qualitative comparisons of interactive 2D and 3D segmentation with SAM, MedSAM and MedicoSAM.
Figure 7: Qualitative results for interactive 2D segmentation. We compare interactive segmentation based on a single point or single box prompt (cyan) with SAM, MedSAM, and MedicoSAM for nine different datasets. For each image, we show prompts with a large improvement of MedicoSAM over SAM and the corresponding MedSAM result. Note that the quantitative results in Fig. 3 show overall improved segmentation with MedicoSAM.
Figure 8: Qualitative results for interactive 3D segmentation. We show rendered 3D segmentations for objects in the volume that were segmented from a single prompt with MedicoSAM.

### v2

@@ Acknowledgment
The work of Anwai Archit was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - PA 4341/2-1. The work of Luca Freckmann was funded by the DFG under Germany’s Excellence Strategy - EXC 2067/1-390729940. This work is supported by the Ministry of Science and Culture of Lower Saxony through funds from the program zukunft.niedersachsen of the Volkswagen Foundation for the ’CAIMed – Lower Saxony Center for Artificial Intelligence and Causal Methods in Medicine’ project (grant no. ZN4257). This work was also supported by the Google Research Scholarship “Vision Foundation Models for Bioimage Segmentation”. We gratefully acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Emmy at NHR@Göttingen as part of the NHR infrastructure, under the project nim00007. We would like to thank Sebastian von Haaren for help with improving manuscript figures and data visualizations.